You need to prepare 200ml of a 0.4M solution of copper sulfate (CuSO4). You have 80mL of a 0.5M CuSO4 solution and 100ml of 7.5% (w/v) CuSO4 solution.

What volume of 7.5% CuSO4 and water needs to be added to the 80mL of 0.5M CuSO4 solution to make up 200mL of 0.4M CuSO4?

Molar mass of CuSO4 = 159.609 g/mol % (weight/volume) means mass (in g) in volume (100mL)
Answer #1

Molarity of 7.5% of CuSO4 solution:

Moles of CuSO4:

7.5 g CuSO4 × ( 1 mol CuSO4 / 159.609 g CuSO4)

= 0.0470 mol CuSO4

Volume of solution:

100 mL × (1 L / 1000 mL) = 0.100 L

Hence molarity = 0.0470 mol / 0.100 L = 0.470 M

Hence molarity and volume of first solution is 0.470 M and 100 mL

Molarity and volume of second solution is 0.5 M and 80 mL

Resultant solution molarity and volume is 0.4 M and 200 mL

M = (M1V1 + M2V2) / (V1+V2)

V1+V2 = 200 mL

M= 0.4 M

M1 = 0.470 M

V1 = ?

M2 = 0.5 M

V2 = 80 mL

0.4 M = (V1 × 0.470 M + 0.5 M × 80 mL) / 200 mL

V1 × 0.470 M = 80 M.mL - 40 M.mL

V1 = 40 M.mL / 0.470 M

V1 = 85.1 mL

Hence the volume of 7.5% solution is 85 mL

amount of water is 200-(80+85) = 35 mL

Hence the volume of 7.5% solution and water is 85+35 = 120 mL